OpenStudy (anonymous):

WRITE 35j^3/10j^2) IN LOWEST TERMS.

OpenStudy (anonymous):

35j^3 7 j^3-2 ------ = -------- 10j^2 2 using \[a ^{m}-a ^{n} = a ^{m-n}\] 7 = --- j 2